OpenStudy (anonymous):

write and explicit formula for the geometric sequence 25,5,1,1/5

OpenStudy (anonymous):

i know An=A*r^n-1

OpenStudy (anonymous):

@whpalmer4

OpenStudy (shamil98):

The ratio is 1/5. r = ratio and you are given the first term (which is A) so \[A_n = 25 * (\frac{ 1 }{ 5 })^{n-1}\]

OpenStudy (anonymous):

wow can you explain that alittle more.. your really smart and still alittle lost @shamil98

OpenStudy (anonymous):

im still alittle lost lol

OpenStudy (shamil98):

The formula is \[A_n = A_1 * r^{n-1}\] Right?

OpenStudy (shamil98):

So you are given the first four terms

OpenStudy (anonymous):

yes i just got it i had to work it out on my calculator. thank you for your help i get it.

OpenStudy (shamil98):

Okay